find a solution to the boundary value problem y²+4y=0 y(π/8)=0, y(π/6)=1 if the general solution to the differential equation is y(x)=c1sin2x + c2cos2x

1
Expert's answer
2022-02-01T16:47:43-0500
"y''+4y=0"

The general solution is


"y=c_1\sin (2x)+c_2 \cos(2x)"

Given "y(\pi/8)=0"


"0=c_1\sin (2(\pi/8))+c_2 \cos(2(\pi/8))"

"c_1+c_2=0"

"c_1=-c_2"

Given "y(\pi/6)=1"


"1=c_1\sin (2(\pi/6))+c_2 \cos(2(\pi/6))"

"c_1(\dfrac{\sqrt{3}}{2})+c_2(\dfrac{1}{2})=1"

"\sqrt{3}c_1+c_2=2"

Then


"c_2=-c_1"

"\sqrt{3}c_1-c_1=2"

"c_1=\dfrac{2}{\sqrt{3}-1}"

"c_1=\sqrt{3}+1"

"c_2=-(\sqrt{3}+1)"

The solution to the boundary value problem is


"y=(\sqrt{3}+1)\sin(2x)-(\sqrt{3}+1)\cos(2x)"

"y=(\sqrt{3}+1)\big(\sin(2x)-\cos (2x)\big)"